# Are Users More Willing to Use Formally Verified Password Managers?

### v1

@@ 5.4.3. Formal Verification Guarantees (RQ3)
As stated before, most scenarios were impactful for users, so in general, our results suggest that participants value the guarantees that formal verification provides. Additionally, we can divide our scenarios into the following categories:
@@
  • Scenarios where a third-party learns all the passwords: scenarios ”Vault Exposure” (S3) and ”Primary Password Exposure” (S4);
@@
  • Scenarios where a third party, over time, can learn a large number of passwords: scenarios ”Unpredictability” (S1), ”Autofill” (S5), and ”Clipboard Clearing” (S6);
@@
  • Scenarios where the user loses access to all the passwords stored in the Password Manager’s vault: scenarios ”Synchronization” (S7) and ”Deleting your vault” (S8);
@@
  • A scenario where the generator does not behave as intended and the passwords generated are not compliant with the users’ inputs: scenario ”Policy Compliance” (S2).
@@
Two of the three most impactful scenarios described a situation where a third party could learn all the users’ passwords (S3 and S4). In these scenarios, 86.5% and 92.5% of participants stated they ”agreed” or ”strongly agreed” that they would stop using a Password Manager if it happened. These scenarios were related to the absence of formal verification where its guarantees would prevent the exposure of passwords to a third party. However, the second, fourth, and fifth most impactful scenarios (S1, S5, and S6) described a situation where a third party, over time, could learn several users’ passwords. In these, over 79% of participants ”agreed” or ”strongly agreed” that they would stop using a Password Manager if they happened. The other type of scenario, where the user loses access to all the passwords they have stored in the Password Manager’s vault (S7 and S8), was not as impactful for participants as the ones mentioned before. Finally, the last type of scenario described a situation where the Password Manager did not work as intended (S2), but its malfunction did not lead to a loss of passwords. It just inconvenienced users. This was the scenario that fewer participants found impactful. Interestingly, the least important feature in our hierarchy—the unpredictability of the password generator (S1)—is a feature that previous work has made efforts to formally verify (grilo22pwdgen, ).
Our results thus suggest that users value the guarantees that formal verification provides. Moreover, users value the guarantees more depending on the impact that said guarantees have. Guaranteeing that passwords are not leaked to third parties (e.g., by ensuring the cryptographic algorithms are well implemented) seems to be very important for participants.
@@
RQ3: Do users value the guarantees that formal verification can provide in Password Managers? Our results suggest that users value the guarantees that formal verification can provide in Password Managers; Our data also seems to indicate that some guarantees are more important than others. For example, guaranteeing that the passwords are not leaked to third parties seems more relevant for participants than guaranteeing that they are synced in the cloud.
6. Discussion
In this section, we situate our results within the broader context of formal verification application. We discuss the main insights and suggest recommendations for future verification efforts and research. Our main insights include:
@@ 6. Discussion
An insight we got was that there is an overall low level of awareness among users about what formal verification is. Despite this, users react positively to formal verification in Password Managers. There appears to be a gap between the technical understanding of formal verification and the perceived value it adds to Password Managers. Some of our participants liked formal verification due to an inherent trust in the mathematical and technical rigor it implies, suggesting that even a superficial understanding or awareness of formal verification may influence user preferences.
@@
A significant insight from this study is that users perceive some Password Manager features as more important. The security of the password vault and the reliability of password generation were highlighted as particularly critical features for users. This insight underscores the need for developers to prioritize formal verification efforts on features that most directly impact user trust and perceived security. Understanding user priorities can guide the allocation of development resources towards the aspects of Password Managers that users value more and enhance the effectiveness and attractiveness of formally verified Password Managers.
6.1. Recommendations
@@
Our results suggest that users do not know what formal verification is. With this in mind, developers of formally verified software should try to convey to users the role of formal verification in their software. We suggest that these efforts should be focused on the specific consequences of formal verification in software instead of theoretical concepts. Avoiding technical language may facilitate the application of best practices such as using metaphors and avoiding jargon (redmiles2017summary, ).
Increasing this transparency may enable users to understand the advantages of formally verified software and, thus, increase its adoption. If formally verified software is valued by users, it may also increase its investment value for businesses (for example, for Password Managers such as Bitwarden (bitwarden, )).
@@
As mentioned before, some users associated formal verification with security (see Section 4.2). While this sometimes can be true, security is not necessarily related to formal verification. As such, we suggest that future efforts to communicate about formal verification should take care to prevent overstating the impact it has. Misleading users can be counterproductive and lead to distrust.
@@
6.2. Limitations
User studies such as these may suffer from bias. Bias can arise from the questions or even the questionnaires. To mitigate our limitations, we performed cognitive interviews [^7] Cognitive interviews involve asking respondents to think aloud as they complete a survey and asking them questions about each survey item (redmiles2017summary, ). to validate the study until no more errors or typos were detected (we did two cognitive interviews for each study). We also followed best practices by offering “don’t know (DK)” or “prefer not to answer” responses (redmiles2017summary, ). Moreover, we asked users to explain their understanding of the topic to mitigate the Dunning-Kruger Effect (e.g., when they stated that they understand what formal verification is, we asked them to explain what it is) and included attention check questions. For example, we include one of these questions among the factors that may impact users’ willingness to use a Password Manager (see Section 5.1) and another among the scenarios in the third section of the survey (see Section 5.1).
We also removed as much jargon as possible. The jargon includes terms like “memory” and “encrypted” but also “formal verification” itself (redmiles2017summary, ) and tried to mitigate the Hawthorne effect [^8] Under the Hawthorne effect users are inclined to agree with researchers (merrett2006reflections, ). by hiding that the study is about formal verification. To prevent bias we randomized the order in which the factors that may impact users’ willingness to use a Password Manager are shown (see Section 5.1), and the order of the scenarios in the third survey section (see Section 5.1).
6.3. Future Work
Exploring formal verification’s impact on users’ willingness to use Password Managers opens up several avenues for further research.
@@ 6.3. Future Work
Our study within the context of Password Managers suggests a positive user reception towards formal verification. Extending this research to other domains where formal verification is applied could offer an insightful research opportunity. For example, investigating user perceptions in the context of autonomous vehicles, medical devices, and blockchain technologies—all of which rely on the integrity and security assurances that formal verification provides—could reveal domain-specific user attitudes and expectations. These studies could help understand whether the positive inclination towards formal verification observed in Password Manager users is universal or if adjustments in communication and implementation strategies are needed based on the application domain.
@@
One significant finding from our study is the general lack of user awareness about formal verification. This gap presents a research opportunity to develop and test new communication methods about formal verification to the public. Inspired by initiatives like the IoT security labels(emami2020ask, ), a “Formal Verification Label” could be designed to provide at-a-glance information about the verification status of a product. Such a label could include simplified symbols or ratings that indicate the extent and areas of formal verification applied, making it easier for users to understand the claims of formally verified software products.
7. Conclusion

### v2

@@ 5.4.3 Formal Verification Guarantees (RQ3)
The scenario’s question also provides insights into whether participants value the guarantees of formal verification in PMs. As stated before, most scenarios were impactful for users, so in general, our results suggest that participants value the guarantees that formal verification provides. We can divide our scenarios into the following categories:
@@
  • Scenarios where a third-party learns all the passwords: “Vault Exposure” (S3) and “Primary Password Exposure” (S4);
@@
  • Scenarios where a third party, over time, can learn manypasswords: “Unpredictability” (S1), “Autofill” (S5), and “Clipboard Clearing” (S6);
@@
  • Scenarios where the user loses access to all the passwords stored in the PM’s vault: “Synchronization” (S7) and “Deleting your vault” (S8);
@@
  • A scenario where the generator does not behave as intended and the passwords are not compliant with the specification: “Policy Compliance” (S2).
Two of the three most impactful scenarios allowed a third party to learn all user passwords (S3, S4). In these scenarios, 86.5% and 92.5% of participants stated they “agreed” or “strongly agreed” that they would stop using a PM if it happened. These scenarios involved the lack of formal verification, which would otherwise prevent password exposure to third parties.
@@
However, the second, fourth, and fifth most impactful scenarios (S1, S5, and S6) described a situation where a third party, over time, could learn several users’ passwords. In these, over 79% of participants “agreed” or “strongly agreed” that they would stop using a PM if they happened. The other type of scenario, where the user loses access to all the passwords they have stored in the PM’s vault (S7 and S8), was not as impactful for participants as the ones mentioned before. Finally, the last type of scenario described a situation where the PM did not work as intended (S2), but its malfunction did not lead to a loss of passwords. It just inconvenienced users. This was the scenario that fewer participants found impactful. Interestingly, the least important feature in our hierarchy, the unpredictability of the password generator (S1), is a feature that previous work has made efforts to formally verify [20].
@@
RQ3: Do users value the guarantees that formal verification can provide in PMs? Our results suggest that users value the guarantees that formal verification can provide in PMs; Our data also seems to indicate that some guarantees are more important than others. For example, guaranteeing that the passwords are not leaked to third parties seems more relevant for participants than guaranteeing that they are synced in the cloud.
6 Discussion
In this section, discuss the main insights and suggest recommendations.
@@ 6 Discussion
An insight we got was that there is an overall low level of awareness among users about what formal verification is. Despite this, users react positively to formal verification in PMs. There appears to be a gap between the technical understanding of formal verification and the perceived value it adds to PMs. Some of our participants liked formal verification due to an inherent trust in the mathematical and technical rigor it implies, suggesting that even a superficial understanding or awareness of formal verification may influence user preferences.
@@
A significant insight from this study is that users perceive some PM features as more important. The security of the password vault and the reliability of password generation were highlighted as particularly critical features for users. This insight underscores the need for developers to prioritize formal verification efforts on features that most directly impact user trust and perceived security. Understanding user priorities can guide the allocation of development resources towards the aspects of PMs that users value more and enhance the effectiveness and attractiveness of formally verified PMs.
6.1 Recommendations
@@
Our results suggest that users do not know what formal verification is. With this in mind, developers of formally verified software should try to convey to users the role of formal verification in their software. We suggest that these efforts should be focused on the specific consequences of formal verification in software instead of theoretical concepts. Avoiding technical language may facilitate the application of best practices such as using metaphors and avoiding jargon [46]. Increasing this transparency may enable users to understand the advantages of formally verified software and, thus, increase its adoption. If formally verified software is valued by users, it may also increase its investment value for businesses (for example, for PMs such as Bitwarden).
@@
As mentioned before, some users associated formal verification with security (see Section˜4.2). While this can sometimes be true, security is not necessarily related to formal verification. As such, we suggest that future efforts to communicate about formal verification should take care to prevent overstating the impact it has. To build trust, communication efforts should carefully distinguish the benefits of verification from broader security properties.
@@
6.2 Limitations
User studies such as these may suffer from bias. Bias can arise from the questions or even the questionnaires. To mitigate our limitations, we performed cognitive interviews [^3] Cognitive interviews involve asking respondents to think aloud as they complete a survey and asking them questions about each survey item [46]. to validate the study until no more errors or typos were detected (we did two cognitive interviews for each study). We also followed best practices by offering “don’t know (DK)” or “prefer not to answer” responses [46]. Moreover, we asked users to explain their understanding of the topic to mitigate the Dunning-Kruger Effect (e.g., when they stated that they understood what formal verification is, we asked them to explain what it is) and included attention check questions. For example, we include one of these questions among the factors that may impact users’ willingness to use a PM (see Section˜5.1) and another among the scenarios in the third section of the survey (see Section˜5.1). We also removed as much jargon as possible. Jargon includes terms like “memory”, “encrypted” and “formal verification” itself [46] and tried to mitigate the Hawthorne effect. To prevent bias, we randomized the order in which the factors that may impact users’ willingness to use a PM are shown (see Section˜5.1), and the order of the scenarios in the third survey section (see Section˜5.1).
6.3 Future Work
Exploring formal verification’s impact on users’ willingness to use PMs opens up several avenues for further research.
@@ 6.3 Future Work
Our study within the context of PMs suggests a positive user reception towards formal verification. Extending this research to other domains where formal verification is applied could offer an insightful research opportunity. For example, investigating user perceptions in the context of autonomous vehicles, medical devices, and blockchain technologies — all of which rely on the integrity and security assurances that formal verification provides — could reveal domain-specific user attitudes and expectations. These studies could help understand whether the positive inclination towards formal verification observed in PM users is universal or if adjustments in communication and implementation strategies are needed based on the application domain.
@@
One significant finding from our study is the general lack of user awareness about formal verification. This gap presents a research opportunity to develop and test new communication methods about formal verification for the public. Inspired by initiatives like the IoT security labels [13], a “Formal Verification Label” could be designed to provide at-a-glance information about the verification status of a product. Such a label could include simplified symbols or ratings that indicate the extent and areas of formal verification applied, making it easier for users to understand the claims of formally verified software products.
7 Conclusion